Grade 12Algebra

Find the angles which are equal to its supplement?????????

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let the measure of the required angle be x0.
We know that, sum of measures of supplementary angle pair is 1800.
Then,
= x + x = 1800 = 2x = 1800
= x = 180/2
= x = 900
Hence, the required angle measures is 900.

Thanks